OpenStudy (anonymous):

Use the division to determine whether the ratios form a proportion. 1/4 and 6/24 A.yes.they form a proportion B.no . they do not form a proportion @cheyennrae1605- I don't if u might know this but u can try ?!

OpenStudy (anonymous):

@cheyennerae1605

Mehek (mehek14):

what do you think? \(\dfrac{6}{24}=\dfrac{1}{4}?\)

OpenStudy (anonymous):

I was thinking yes but not sure cause I just started learning bout proportions

Mehek (mehek14):

if you divide \(\dfrac{6}{24}\) by 6 do you get \(\dfrac{1}{4}\)?

OpenStudy (anonymous):

all I see is 0.0416666666?

Mehek (mehek14):

what is 6/6?

OpenStudy (anonymous):

1

Mehek (mehek14):

24/6?

OpenStudy (anonymous):

4

Mehek (mehek14):

so you get \(\dfrac{1}{4}\)

Mehek (mehek14):

which means they are proportional

OpenStudy (anonymous):

oh ok I get it thx so much !

Mehek (mehek14):

yw \(\ddot\smile\)